# Out-of-Distribution Detection in Molecular Complexes via Diffusion Models for Irregular Graphs

### v1

@@ Abstract
Predictive machine learning models generally excel on in-distribution data, but their performance degrades on out-of-distribution (OOD) inputs. Reliable deployment therefore requires robust OOD detection, yet this is particularly challenging for irregular 3D graphs that combine continuous geometry with categorical identities and are unordered by construction. Here, we present a probabilistic OOD detection framework for complex 3D graph data built on a diffusion model that learns a density of the training distribution in a fully unsupervised manner. A key ingredient we introduce is a unified continuous diffusion over both 3D coordinates and discrete features: categorical identities are embedded in a continuous space and trained with cross-entropy, while the corresponding diffusion score is obtained analytically via posterior-mean interpolation from predicted class probabilities. This yields a single self-consistent probability-flow ODE (PF-ODE) that produces per-sample log-likelihoods, providing a principled typicality score for distribution shift. We validate the approach on protein-ligand complexes and construct strict OOD datasets by withholding entire protein families from training. PF-ODE likelihoods identify held-out families as OOD and correlate strongly with prediction errors of an independent binding-affinity model (GEMS), enabling a priori reliability estimates on new complexes. Beyond scalar likelihoods, we show that multi-scale PF-ODE trajectory statistics–including path tortuosity, flow stiffness, and vector-field instability–provide complementary OOD information. Modeling the joint distribution of these trajectory features yields a practical, high-sensitivity detector that improves separation over likelihood-only baselines, offering a label-free OOD quantification workflow for geometric deep learning.
@@ 1 Introduction
Out-of-distribution (OOD) detection is essential in machine learning, because real-world data often does not match the data distribution used for model training. In these cases, models often make overly confident predictions on the unfamiliar or shifted inputs, which can lead to serious errors, especially in high-stakes domains like diagnostics or drug discovery. Effective OOD detection provides a safeguard by identifying when a model is operating outside its reliable domain, allowing users to reject predictions or handle uncertain cases more appropriately. This not only improves safety but also builds trust in machine learning models deployed in dynamic environments.
Here we demonstrate that denoising diffusion probabilistic models (DDPMs) can serve as practical OOD detectors that can be integrated with deployed predictive machine learning models to provide a supporting confidence score and flag predictions made on input data that falls outside the model’s reliable domain. These models can be trained on any data amendable to continuous diffusion and require no labels. When trained on the same training set as a predictive machine learning model, the DDPM learns a probability density function of the training distribution. For new graphs, it can output a scalar likelihood score that indicates the degree to which new data samples match the training distribution. If presented with an OOD data point, the DDPM produces a low likelihood value, signaling that the input data point is atypical under the training distribution and that the predictive machine learning model’s output may be unreliable.
To estimate this OOD likelihood, we make use of the specific properties of DDPMs: These generative models define a continuous forward process that gradually transforms a clean input structure into pure Gaussian noise using a stochastic differential equation (SDE). Then, they learn a corresponding reverse process to travel back from noise to the original clean structure by solving the backward SDE. Importantly, the SDE that governs the noising process has a corresponding, mathematically equivalent probability-flow ordinary differential equation (PF-ODE). This deterministic ODE transports samples from the noisy state back to the clean data state while matching the probability distribution of the SDE at every intermediate time point. In other words, the ODE transforms the simple Gaussian distribution of the noise into the complex distribution of the real data. In this setting, we can estimate for any new data point how well it matches the training distribution of the diffusion model by integrating the divergence of the probability flow ODE’s drift term along the trajectory from the clean state to the terminal noisy state. The resulting log-likelihood value provides a mathematically rigorous estimator of how out-of-distribution the new data is.
We showcase the OOD detection ability of DDPMs using a database of 3D protein-ligand interaction complexes called PDBbind [1, 2]. This data combines discrete attributes (atom and residue types) with continuous spatial geometry (atom and residue positions). Therefore, these interactions are typically modeled as complex unordered graphs, making OOD detection with simple statistical checks or similarity computations impractical. Instead, we need methods that produce OOD scores based on a joint assessment of discrete and geometric graph features. DDPMs can accomplish this task, because they explicitly model the probability density of this joint distribution and are amenable to log-likelihood-based evaluation.
@@
To train a diffusion model on this complex data and obtain a probability flow ODE for log-likelihood computation, we introduce a unified continuous diffusion framework that treats both 3D geometry and discrete chemical identities in a single Euclidean state. By mapping the categorical data onto a continuous, spherical embedding space and concatenating these chemical embeddings with the 3D spatial coordinates, we can subject the entire molecular state to a single diffusion process. A single SE(3)-equivariant graph neural network (EGNN) consumes this joint state of coordinates and noisy categorical embeddings and outputs per-node clean coordinates and probability (logits) of each chemical type. We then calculate a posterior mean (essentially a weighted average of the learned chemical prototypes based on the predicted probabilities) and compute the score analytically from this ”clean” signal. This method of posterior-mean interpolation allows us to train the model using standard cross-entropy loss while maintaining the smooth, continuous dynamics required for the diffusion process. After training, the model defines a single, self-consistent probability flow ODE whose pathwise divergence integrated in time, together with a terminal Gaussian evaluation, yields per-complex log-likelihoods.
Trained on a distribution of known protein-ligand interactions, our diffusion model allows to sample new interactions with realistic molecular structure, demonstrating successful distribution learning (Section S1.6). Crucially, it allows to estimate the in-distribution (ID) and out-of-distribution (OOD) likelihood for new interactions using its probability flow ODE. To demonstrate this ability, we constructed a custom train-test split of the PDBbind database by isolating seven protein families from the training data, thereby forming seven strict OOD test sets. We combined these test sets with two datasets exhibiting intermediate degrees of distribution shift and validated the OOD degrees of these datasets in an extensive bioinformatic analysis. For eight out of the nine datasets, the diffusion log-likelihood distributions correctly identified the dataset’s OOD level relative to the training data.
Furthermore, we hypothesized that OOD complexes induce atypical diffusion trajectories, capturing the model’s struggle to process OOD structures. Therefore, we enriched the log-likelihood signal with 18 additional descriptors extracted along the diffusion trajectory, including path tortuosity, flow stiffness, vector-field instability and energetic cost. These quantities act as trajectory-level fingerprints of how “hard” the model has to work to explain a given complex. When modeled jointly with log-likelihoods, this augmented feature space yields a substantially stronger OOD detector, which consistently distinguishes ID from OOD complexes. Notably, this augmented approach correctly classified the dataset that had previously been misidentified based on log-likelihoods alone, showing that OOD structures induce distinct dynamics in their diffusion trajectories that facilitate OOD detection.
As a prove of usability in combination with a predictive model, we train the deep-learning-based protein-ligand binding affinity prediction method GEMS [3] on the same data. GEMS showed significantly reduced accuracy on test datasets that had been flagged as OOD by our diffusion model. This demonstrates that diffusion-based OOD estimates can provide an a priori assessments of GEMS’ reliability and accuracy on given new data, based on the degree of representation of the new data in the training distribution of both models.
A central finding of this work is that, despite working in a completely unsupervised setting, the diffusion model still recovers chemically meaningful structure. The log-likelihoods assigned by its probability-flow ODE correctly rank held-out protein families as increasingly out-of-distribution, align with bioinformatically computed protein–ligand similarity scores, and correlate with the prediction errors of deterministic pre-existing models across datasets. This agreement is non-trivial: it shows that a model which only sees raw 3D interaction geometry and discrete types can reconstruct many of the same distinctions that emerge from strongly biased, affinity-supervised or hand-engineered scores. As a result, trajectory statistics collected along the probabilty-flow ODE provide an OOD signal that is both task-agnostic and chemically interpretable, and can serve as an independent check on the effective training distribution seen by downstream predictive models.
@@ 2 Results
A practical obstacle in structure-based machine learning is that labels are scarce, but structures are not. As of December 2025, the Protein Data Bank (PDB) contains more than $2.4\times 10^{5}$ released, atomic-resolution structures, with on the order of $10^{4}$ new structures added per year [4]. By contrast, widely used structure–affinity datasets are orders of magnitude smaller: PDBbind v2020 provides 19,44319{,}443 protein-ligand complexes with curated binding affinities, and evaluation benchmarks such as CASF-2016 comprise only a few hundred labeled co-crystal complexes [5, 6].
This label-free regime is precisely where our OOD detection approach is designed to operate. Our diffusion model is trained only to model the joint distribution of 3D geometry and discrete atom/residue identities. It never uses binding affinities, protein-family annotations, hand-crafted similarity scores, or any OOD labels. Nevertheless, PF-ODE log-likelihoods and trajectory statistics provide a quantitative, per-complex notion of typicality with respect to the training distribution. In the next sections we show that this label-free signal identifies held-out protein families as OOD and anticipates the error profile of an independent affinity predictor trained with supervision.
To evaluate the diffusion model’s OOD detection capability, we used a custom data split of the PDBbind (v.2020) dataset following the procedure of [7]. We constructed seven OOD test sets by isolating entire protein families from the dataset. The remaining data was subjected to CleanSplit filtering [3] and randomly split into training and validation sets. The filtering allowed us to include CASF2016 as an additional test set without data leakage. This resulted in a total of nine non-training datasets representing a spectrum of distribution shifts relative to the training data: The validation set, drawn from the same distribution, should exhibit minimal shift. The CASF2016 dataset occupies an intermediate position. While originally sampled from the training data, CleanSplit filtering removed structural overlaps, likely inducing a moderate shift. The largest shifts are expected for the datasets of isolated protein families, denoted by their cluster centers 1nvq, 1sqa, 2p15, 2vw5, 3dd0, 3f3e, and 3o9i. Following training of the diffusion model, we computed log-likelihoods and 18 additional PF-ODE trajectory statistics across all datasets and evaluated the discriminative power of these features in detecting OOD inputs.
@@ 2.1 Dataset Log-Likelihood Distributions
The log-likelihoods assigned to the training, validation, and OOD datasets generally aligned with expectations (Figure 1). As anticipated, the training dataset received high log-likelihoods, showing that this dataset fits the distribution learned by the diffusion model. The validation dataset received log-likelihoods slightly lower than the training set, which is expected for unseen samples drawn from the training distribution. The OOD datasets composed of extracted protein family clusters mostly yielded low log-likelihoods, indicating they were correctly identified as OOD. Finally, the CASF2016 dataset produced log-likelihoods intermediate to the training and OOD datasets. This aligns with the nature of the CleanSplit filtering used to ensure the independence of CASF2016: While it removes high-similarity complexes, it does not enforce the strict independence achieved by holding out entire protein families. Among the OOD datasets, the HIV protease family showed the most pronounced distribution shift in log-likelihoods relative to the training data. This finding aligns with biological expectations, as HIV protease complexes possess unique characteristics. These viral proteases contain an active site located between two short identical subunits, flanked by flexible molecular flaps that close upon substrate binding. Inhibitors targeting these enzymes are typically large molecules designed to mimic the proteolytic transition state without being cleaved, effectively blocking the active site. A significant deviation from biological expectation occurred in one of the nine datasets. The carbonic anhydrases in the 3dd0 cluster yielded high log-likelihoods, suggesting that this dataset is in-distribution relative to the training data. We focus our evaluation of the PF-ODE log-likelihoods on the remaining eight datasets and discuss the 3dd0 results separately in Sections 2.3, S1.1 and S1.4.
Log-Likelihood Variance
The distributions of log-likelihoods (Figure 1) were broad across all datasets. This variance is expected for the inherently diverse training and validation datasets, but also for CASF2016 dataset, which is diverse by design. OOD datasets generally showed lower variance in log-likelihoods. Nevertheless, a considerable spread remained also in the OOD datasets, suggesting that these datasets still contain typical as well as atypical complexes. This is likely attributable to our splitting strategy: While we sequestered specific protein families, the ligand space remained uncontrolled. Consequently, OOD pockets may bind a diverse mix of ligands, ranging from familiar to exotic structures, which naturally drives variation in the likelihood scores. The impact of ligand diversity on log-likelihood spread becomes evident by analyzing the ligand size distribution and chemical diversity within the datasets (Figures S3 and S6). For every dataset, we identified a medoid ligand, defined as the molecule with the smallest sum of distances (1Tanimoto similarity1-$\text${Tanimoto similarity}) to all other ligands, and computed the distance of all other ligands to this central reference. We found a strong correlation between cluster tightness (median distance to the medoid) and the standard deviation of the log-likelihood distributions (Pearson’s $r=0.832$). This shows that datasets with diverse sets of ligands exhibited the broadest log-likelihood distributions. The most extreme case is the cluster of transporter protein binding pockets 3f3e, which exhibits a ligand size distribution much broader than all other datasets (Figure S3).
@@
Figure 1: Out-of-distribution datasets yield lower log-likelihoods: Distributions of log-likelihoods assigned to protein-ligand complexes belonging to the a) training, b) validation, c) CASF2016 and the out-of-distribution (OOD) datasets d) 1nvq, e) 1sqa, f) 2p15, g) 2vw5, h) 3f3e and i) 3o9i. Lower values signify increased deviation from the learned distribution (OOD) and higher values indicate high in-distribution (ID) probability. Means, medians, standard deviations (Std) and number of samples (N) are depicted on the right side of the histograms. Vertical dashed lines indicate the median of the respective log-likelihood distributions. Distributions were subjected to individual outlier removal using the IQR method (1.5×IQR rule) followed by min-max normalization to a [-1, 0] range using global minimum and maximum values across all distributions. The training distribution was randomly subsampled, showing only a fraction of all 10’510 complexes.
2.2 Bioinformatic Validation of OOD Log-Likelihoods
To validate the OOD log-likelihoods obtained from the diffusion model, we performed an extensive bioinformatic analysis, evaluating the degree of representation of the validation, CASF2016 and OOD datasets in the training distribution. First, we computed pairwise similarities across all $19^{\prime}443$ complexes in PDBbind using a combined assessment of protein similarity (TM-align [8]), ligand similarity (Tanimoto scores [9]) and binding conformation similarity (pocket-aligned ligand RMSD). Tanimoto similarity is commonly used to measure the similarity between small molecules. Based on comparing chemical fingerprints, this score ranges between 0 (no similarity) and 1 (identical) and identifies compounds with similar structural and chemical properties. TM-align compares protein structures by finding the optimal alignment of their three-dimensional shapes, producing a TM-score ranging from 0 (very dissimilar fold) to 1 (identical fold). The pocket-aligned ligand RMSD calculation completes the similarity assessment by comparing ligand positioning within aligned protein pockets (based on TM-align). For each metric, we computed all $189*10^{6}$ pairwise similarities across PDBbind and generated individual pairwise similarity matrices. Combining these three metrics allows for a robust structure-based similarity quantification between protein-ligand complexes.
We approximated the degree of representation of the non-training datasets within the training data by quantifying the number and the magnitude of similarities shared between the training complexes and the respective non-training datasets. We then aligned these similarity results with the log-likelihood distributions obtained from the diffusion model (Figure 2).
  • TM scores aligned closely with the log-likelihood distributions (Figure 2a, c). Datasets sharing fewer and weaker similarities in protein folds with the training data exhibit lower log-likelihoods (suggesting more OOD). While validation set proteins were well-represented within the training data, the CASF2016 proteins shared only moderate structural similarity. The proteins of the OOD datasets were highly underrepresented in the training data, confirming that the pocket-level splitting mechanism effectively isolates proteins with distinct folds (Figure 2c).
  • Tanimoto scores showed a slightly weaker correlation with the log-likelihood distributions (Figure 2a, b). Nevertheless, a similar trend persisted: Datasets sharing weaker ligand similarities with the training set yield lower log-likelihoods. The OOD datasets showed significantly lower similarity to the training data than the validation set. This confirms that isolating specific protein types during OOD splitting automatically induces a moderate shift in the ligand distribution.
  • Aggregated similarity scores were derived by computing the sum of Tanimoto, TM-align, and inverted RMSD scores for each pair of complexes. The aggregated similarity scores between two datasets offer a more complete measure of independence, as it accounts for protein, ligand and binding pose similarity simultaneously. The number and magnitude of shared similarities aligned well with the log-likelihood distributions (Figure 2a, d). Datasets with complexes sharing fewer and weaker similarities with the training data obtained lower log-likelihoods.
While the magnitude of Tanimoto and TM scores shared between training and non-training datasets certainly influences the log-likelihood distributions, these metrics should not be taken as a direct proxy for the number of similar complexes shared between them. This is because similar protein folds may be associated with entirely different ligands, and structurally similar ligands might bind to completely disparate proteins, thereby resulting in the absence of similarity at the protein–ligand interaction level. Only the combination of these metrics at the complex level, combined with a supplementary assessment of binding conformation similarity, can accurately quantify the structural overlap between two datasets. This combined assessment demonstrates that all designed OOD datasets share fewer and weaker complex-level similarities with the training data: For many training complexes, no representation of the same binding pattern is found in these datasets whatsoever (aggregated similarity scores at zero). The similarity scores of the CASF2016 dataset confirm that the CleanSplit filtering applied to the training data renders the CASF2016 dataset more independent, without reaching the same strict level of independence achieved by the OOD benchmarks.

### v2

@@ Abstract
Predictive machine learning models generally excel on in-distribution data, but their performance degrades on out-of-distribution (OOD) inputs. Reliable deployment of machine learning models therefore requires robust OOD detection, yet identifying OOD data in irregular 3D graphs remains a fundamental challenge. We present the first unsupervised OOD detection framework specifically designed for 3D geometric graphs, using a generative diffusion model to approximate the joint distribution of geometry and discrete features. We demonstrate that probability-flow ODE trajectories provide a quantitative, per-complex notion of typicality with respect to the training distribution. By augmenting likelihoods with eighteen trajectory-level geometric features we effectively bypass the complexity bias inherent in standard density estimation. Our framework accurately identifies OOD protein-ligand complexes and anticipates the error profiles of independent binding-affinity predictors without requiring task-specific labels. In this sense, this OOD quantification can act as a certificate for benchmark datasets and the connected generalization claims.
Figure 1: Schematic of out-of-distribution (OOD) detection via PF-ODE trajectory analysis: a) Stochastic Differential Equation (SDE): The diffusion framework defines a forward process that transforms the input data distribution (protein-ligand complexes) into a Gaussian noise distribution via stochastic, noisy paths. b) Probability Flow ODE (PF-ODE): The trained model learns a deterministic vector field representing the average marginal flow of the SDE. This defines a continuous, reversible mapping between the noise and data spaces, transporting every sample along a unique, deterministic path (trajectory). c) Comparison of trajectories for in-distribution (ID) versus OOD samples: ID samples follow efficient, direct paths through regions where the vector field is well-constrained by training data. In contrast, OOD samples traverse low-density regions where the learned flow is ambiguous and chaotic due to a lack of training samples, resulting in erratic, high-tortuosity paths. Geometric features extracted from these trajectories, such as path efficiency, flow energy, and Lipschitz stability, serve as robust signals to detect OOD inputs.
@@ 1 Introduction
Out-of-distribution (OOD) detection is essential in machine learning, because real-world data often does not match the data distribution used for model training. In these cases, models often make overly confident predictions on unfamiliar or shifted inputs, which can lead to serious errors, especially in high-stakes domains like diagnostics or drug discovery. Effective OOD detection provides a safeguard by identifying when a model is operating outside its reliable domain, allowing users to reject predictions or handle uncertain cases more appropriately. This not only improves safety but also builds trust in machine learning models deployed in dynamic environments.
Traditional OOD detection approaches can be broadly categorized into supervised discriminative methods and unsupervised generative methods (see Section S1.1). Discriminative approaches typically leverage scores derived from pretrained classifier outputs [15, 22, 24, 34]. However, their dependence on in-distribution labels limits their applicability in unsupervised settings where such labels are unavailable. Consequently, unsupervised approaches typically employ generative models and fall into two primary sub-categories: reconstruction-based and likelihood-based methods.
Reconstruction-based approaches [2, 37, 3] operate on the assumption that a generative model trained on in-distribution data will fail to accurately reconstruct OOD samples. However, these methods frequently suffer from reconstruction shortcuts, where powerful decoders successfully reconstruct simple OOD inputs despite their semantic novelty, rendering the reconstruction error metric unreliable [5, 41].
Alternatively, likelihood-based methods explicitly model the probability density of the training data, flagging test samples in low-density regions as OOD. These approaches frequently employ denoising diffusion probabilistic models (DDPMs) to capture this density. By learning the probability density function of the training distribution, the DDPM can output a scalar likelihood score indicating the degree to which a new sample matches the learned distribution [32, 12]. While theoretically sound, these models often struggle with the complexity bias [27, 30, 19], where low-complexity OOD data is assigned high likelihoods, leading to false negatives. To mitigate these limitations, Heng et al. (2024) [16] demonstrated that the rate-of-change and curvature of diffusion trajectories can serve as effective OOD signals for image benchmarks. Their trajectory-focused approach proved resistant to the complexity bias that often confounds likelihood-based OOD classification.
Crucially, while OOD detection has been adapted for graph data [36, 26, 21, 31], existing methods predominantly focus on 2D topological graphs [26, 21, 31]. These approaches neglect the precise 3D geometry that is essential for modeling 3-dimensional data such as molecular complexes, leaving a critical gap in the growing field of geometric deep learning.
Here, we present a novel OOD detection framework built upon a diffusion model tailored for 3D irregular graphs. By modeling likelihood scores jointly with an expanded set of 18 distinct diffusion trajectory features, we derive complementary, robust OOD detection signals. This strategy effectively bypasses the complexity bias inherent in isolated log-likelihood evaluation, enabling robust OOD detection within the 3D graph domain. We showcase this OOD detection approach using a database of protein-ligand interaction complexes called PDBbind [35, 25], which are typically modeled as 3-dimensional geometric graphs combining discrete attributes (atom and residue types) with continuous spatial coordinates. Despite the complexity of this domain, our framework accurately identifies OOD complexes characterized by protein pockets unseen during training.
@@
To train a diffusion model on this complex data that combines 3D geometry and discrete graph features, we introduce a unified continuous diffusion framework that unifies both 3D geometry and discrete features within a single Euclidean state. By mapping the categorical data onto a continuous, spherical embedding space and concatenating these chemical embeddings with the 3D spatial coordinates, we can subject the entire molecular state to a single diffusion process. A single SE(3)-equivariant graph neural network (EGNN) consumes this joint state of coordinates and noisy categorical embeddings and outputs per-node clean coordinates and probability (logits) of each chemical type. We then calculate a posterior mean (essentially a weighted average of the learned chemical prototypes based on the predicted probabilities) and compute the score analytically from this ”clean” signal. This method of posterior-mean interpolation allows us to train the model using standard cross-entropy loss while maintaining the smooth, continuous dynamics required for the diffusion process.
To detect OOD samples using this diffusion model trained on in-distribution protein-ligand complexes, we make use of the specific properties of DDPMs: These generative models define a continuous forward process that gradually transforms a clean input structure into pure Gaussian noise using a stochastic differential equation (SDE, Figure 1a). Then, they learn a corresponding reverse process to travel back from noise to the original clean structure by solving the backward SDE. This admits a mathematically equivalent formulation of the diffusion process known as the probability-flow ordinary differential equation (PF-ODE) [32]. This ODE defines a deterministic vector field representing the average marginal flow of the SDE, establishing a continuous, reversible mapping between the noise and data spaces (Figure 1b). This mapping transports every sample along a unique, deterministic path, which we refer to as its PF-ODE trajectory (Figure 1c).
In this setting, we can compute the exact log-likelihood of any sample by integrating the divergence of the drift term along its PF-ODE trajectory. While this score theoretically estimates how OOD a sample is, it has been repeatedly shown to suffer from complexity bias, where very simple inputs are often assigned in-distribution likelihoods despite being OOD [27, 30, 19]. To mitigate this, we augment the likelihood scores with additional features extracted from the sample’s PF-ODE trajectories (Figure 1c). The underlying idea is that in-distribution (ID) and OOD samples are distinguishable by the geometric characteristics of their PF-ODE trajectories: ID samples follow efficient, direct paths through regions where the vector field is well-constrained by training data. In contrast, OOD samples traverse low-density regions where the learned flow is ill-defined due to a lack of similar training samples, resulting in erratic and chaotic trajectories. Features extracted from these trajectories, such as path efficiency, flow energy, and Lipschitz stability, serve as robust signals to detect OOD inputs.
To demonstrate the efficacy of this trajectory-focused approach in the 3D graph domain, we constructed a custom train-test split of the PDBbind database. We formed seven strict OOD test sets by systematically excluding entire protein families from the training data. This rigorous exclusion ensures that these proteins are never encountered during training and appear novel and atypical to the model at inference. Additionally, we incorporated the CASF2016 benchmark set and a standard validation set to serve as intermediate level OOD test sets.
After training the diffusion model on this specific split, we quantified the OOD signal inherent in PF-ODE trajectories and tested its ability to flag the OOD test sets as anomalies. Our analysis demonstrates that the derived log-likelihoods contain considerable OOD signal and are chemically meaningful. Across eight out of nine datasets, they correctly rank held out protein families as increasingly OOD, aligning with bioinformatically computed similarities. These likelihoods even correlate with the prediction errors of an independent binding affinity prediction model that was trained on the same data. However, despite this alignment, the log-likelihood distributions retained significant overlap with the training data, making robust separation of OOD and ID samples difficult. Crucially, for one dataset comprising low complexity interactions, the likelihoods wrongly suggested ID status, which indicates a severe complexity bias. Concluding that log-likelihoods alone are insufficient for reliable OOD detection, we incorporated additional PF-ODE trajectory statistics. This augmented feature space yields a substantially stronger OOD detector that consistently distinguishes ID from OOD complexes, successfully classifying the low complexity dataset that was previously misidentified.
Overall, our results indicate that diffusion models can serve as practical OOD detectors. They can be integrated with deployed predictive machine learning models to provide a supporting confidence score and flag unreliable predictions. Crucially, we show that augmenting log-likelihoods with geometric features of the diffusion trajectories significantly increases detection accuracy and provides immunity against the complexity bias. While validated here within the challenging 3D graph domain, this OOD detection framework is inherently generalizable to any data modality amenable to continuous diffusion and requires no labels.
@@ 2 Results
To rigorously evaluate the diffusion model’s OOD detection capability, we used a custom data split of the PDBbind (v.2020) dataset [20] with nine test datasets designed to simulate a spectrum of distribution shifts relative to the training data.
  • Minimal shift (validation set): A standard validation set drawn from the training distribution.
  • Intermediate shift (CASF2016): We applied CleanSplit filtering [13] to the training data, removing structural overlaps with the CASF-2016 benchmark.
  • Strong shift (isolated protein families): We constructed seven strict OOD sets by isolating entire protein families from the training data, including serine/threonine-protein kinases, estrogen receptors, HIV proteases, $\alpha$-carbonic anhydrases, urokinase-type plasminogen activators, HSP82 chaperones, and a cluster of transporter proteins (cluster centers 1nvq, 2p15, 3o9i, 3dd0, 1sqa, 2vw5, 3f3e; see Section 4.1).
In the following sections, we first validate the OOD level of our test sets through an extensive bioinformatic similarity analysis, and then analyze the power of the diffusion model to detect them. We evaluate OOD detection performance based on PF-ODE log-likelihoods alone, as well as in combination with 18 additional trajectory statistics.
2.1 Bioinformatic Validation of OOD Levels
To confirm the OOD levels of our test datasets, we analyzed the extent to which the validation, CASF2016, and the isolated OOD datasets are represented within the training data. We approximated this degree of representation by quantifying the number and the magnitude of similarities shared between the training complexes and the respective test dataset (Figure 2). We computed pairwise similarities across all $19^{\prime}443$ complexes in PDBbind using a combined assessment of protein similarity, ligand similarity and binding conformation similarity. TM-align [40] compares protein structures by finding the optimal alignment of their three-dimensional shapes, producing a TM-score ranging from 0 (very dissimilar fold) to 1 (identical fold). Tanimoto similarity [4] is commonly used to measure the similarity between small molecules. Based on comparing chemical fingerprints, this score ranges between 0 (no similarity) and 1 (identical) and identifies compounds with similar structural and chemical properties. The pocket-aligned ligand root mean squared deviation (RMSD) completes the similarity assessment by comparing ligand positioning within aligned protein pockets.
The distributions of TM-scores (Figure 2b) reveal that proteins in the OOD datasets are severely underrepresented in the training data, confirming that the pocket-level splitting mechanism effectively isolates proteins with distinct folds. As anticipated, the validation set proteins are well represented withing the training data, while CASF-2016 proteins exhibit moderate similarity. Regarding ligand similarities, Tanimoto distributions (Figure 2a) indicate that the OOD dataset ligands share lower similarity with the training data ligands than the validation set. This confirms that isolating specific protein types during our data splitting automatically induces a moderate shift in the ligand distribution.
While the Tanimoto and TM scores alone are informative, they are insufficient proxies for protein-ligand interaction similarity. This is because similar protein folds may be associated with entirely different ligands, and structurally similar ligands might bind to completely disparate proteins, thereby resulting in the absence of similarity at the protein–ligand interaction level. Only the combination of these metrics at the complex level, combined with a supplementary assessment of binding conformation similarity, can accurately quantify the structural similarity of protein-ligand complexes. This information is captured by the aggregated similarity scores (Figure 2c), calculated as the sum of Tanimoto, TM-align, and inverse ligand RMSD for each pair. This combined metric demonstrates that all OOD datasets share fewer and weaker complex-level similarities with the training data than the validation set. For many training complexes, no representation of the same binding pattern is found in these datasets whatsoever (aggregated similarity scores at zero). This confirms that these datasets are OOD with respect to the training data. The CASF2016 scores confirm that this dataset is moderately OOD and does not reach the same strict level of independence achieved by the OOD datasets.
Figure 2: Bioinformatic similarity analysis confirms OOD levels: Comparison of the distributions of train-test similarity scores. a) Ligand similarity: Calculated using Tanimoto scores between count-based molecular fingerprints. High scores indicate high similarity (1.0 = identical). b) Protein similarity: Determined by TM-align based on optimal 3D protein structure alignment. High scores indicate high structural similarity (1.0 = identical) c) Aggregated similarity: A composite score S=max(Tanimoto+TMScore+(1RMSD),0)S=max(Tanimoto+TMScore+(1-RMSD),0) was calculated as the sum of Tanimoto similarity, TM-scores, and inverted pocket-aligned ligand root mean squared deviation (RMSD). High scores signify highly similar protein-ligand complexes (3.0 = identical). Each box represents the distribution of $N=10,510$ similarity scores. Specifically, these are the scores between each training dataset complex and its most similar counterpart in the respective test dataset, as measured by that specific metric. Boxplots show the median (centre line), 25th–75th percentiles (box), whiskers extend to data points within 1.5 × IQR, outliers are not shown.
@@ 2.2 Dataset Log-Likelihood Distributions
The log-likelihoods assigned to the training, validation, and OOD datasets generally aligned with expectations (Figure 3). As anticipated, the training dataset received high log-likelihoods, showing that this dataset fits the distribution learned by the diffusion model. The validation dataset received log-likelihoods slightly lower than the training set, which is expected for unseen samples drawn from the training distribution. The OOD datasets composed of extracted protein family clusters mostly yielded low log-likelihoods, indicating they were correctly identified as OOD. Finally, the CASF2016 dataset produced log-likelihoods intermediate to the training and OOD datasets. This aligns with the nature of the CleanSplit filtering used to ensure the independence of CASF2016: While it removes high-similarity complexes, it does not enforce the strict independence achieved by holding out entire protein families. Among the OOD datasets, the HIV protease family showed the most pronounced distribution shift in log-likelihoods relative to the training data. This finding aligns with biological expectations, as HIV protease complexes possess unique characteristics. These viral proteases contain an active site located between two short identical subunits, flanked by flexible molecular flaps that close upon substrate binding. Inhibitors targeting these enzymes are typically large molecules designed to mimic the proteolytic transition state without being cleaved, effectively blocking the active site.
@@
Figure 3: Out-of-distribution datasets yield lower log-likelihoods: Distributions of log-likelihoods assigned to protein-ligand complexes belonging to the a) training, b) validation, c) CASF2016 and the out-of-distribution (OOD) datasets d) 1nvq, e) 1sqa, f) 2p15, g) 2vw5, h) 3dd0, i) 3f3e and j) 3o9i. Lower values signify increased deviation from the learned distribution (OOD) and higher values indicate high in-distribution (ID) probability. Means, medians, standard deviations (Std) and number of samples (N) are depicted on the right side of the histograms. Vertical dashed lines indicate the median of the respective log-likelihood distributions. Distributions were subjected to individual outlier removal using the IQR method (1.5×IQR rule) followed by min-max normalization to a [-1, 0] range using global minimum and maximum values across all distributions. The training distribution was randomly subsampled, showing only a fraction of all 10’510 complexes.
The spread of the log-likelihoods distributions was notably large across all datasets. This is expected for the inherently diverse training and validation datasets, but also for CASF2016, which is diverse by design. While the OOD datasets of extracted protein families generally showed lower log-likelihood variance, a considerable spread remained, suggesting that these datasets still contain typical as well as atypical complexes. This is likely attributable to our splitting strategy: While we sequestered specific protein families, the ligand space remained uncontrolled. Consequently, OOD pockets may bind a diverse mix of ligands, ranging from familiar to exotic structures, which naturally drives variation in the likelihood scores. The impact of ligand diversity on log-likelihood spread becomes evident by analyzing the ligand size distribution and chemical diversity within the datasets (Figures S2 and S3). For every dataset, we identified a medoid ligand, defined as the molecule with the smallest sum of distances (1Tanimoto similarity1-$\text${Tanimoto similarity}) to all other ligands, and computed the distance of all other ligands to this central reference. We found a strong correlation between cluster tightness (median distance to the medoid) and the standard deviation of the log-likelihood distributions (Pearson’s $r=0.832$). This shows that datasets with diverse sets of ligands exhibited the broadest log-likelihood distributions. The most extreme case is the cluster of transporter protein binding pockets 3f3e, which exhibits a ligand size distribution much broader than all other datasets (Figure S2).
A significant deviation from biological expectation was observed in the log-likelihood distribution of the $\alpha$-carbonic anhydrase dataset (3dd0, Figure 3h). This dataset yielded unexpectedly high log-likelihoods, suggesting that it is more in-distribution than the training data itself. We attribute this phenomenon to the complexity bias inherent in deep generative models, where high probability densities are often assigned to simple inputs regardless of their semantic novelty [27]. The 3dd0 complexes exhibit markedly lower structural complexity (e.g., fewer ligand atoms) compared to the training dataset and all other OOD datasets (Figure S2). Consequently, we focus the following evaluation of PF-ODE log-likelihoods on the remaining eight datasets and reserve the discussion of 3dd0 and its successful detection via trajectory features for Section 2.5.
2.3 Log-Likelihoods Align with Bioinformatic OOD Levels
We observed a strong alignment between the diffusion model’s log-likelihood distributions and the bioinformatic OOD levels (Figure 4). These bioinformatic scores are distributions of similarity scores between dataset complexes, calculated as the sum of Tanimoto, TM-align, and inverted RMSD scores to account for protein, ligand, and binding pose simultaneously (Section 2.1). These results demonstrate that datasets sharing fewer and weaker complex-level similarities with the training data generally yield lower log-likelihoods. This direct correspondence confirms that the model’s likelihood estimates are chemically grounded, capturing the structural novelty of the underlying molecular interactions.